OpenStudy (anonymous):

What is the volume of 320.0 g of mercury vapor at 152°C and 0.775 atm?

OpenStudy (anonymous):

Without a compressibility factor for mercury, let's assume mercury vapor is an ideal gas\[PV = {m \over MM} RT \rightarrow V = {m \cdot R \cdot T \over MM \cdot P}\] Be sure to use the correct gas constant such that the units are correct. 8.314 is valid if you convert Celcius to Kelvin. You will get volume units of Liters.